Upon complete combustion of 172 g of an organic substance containing C, H, O and having mass
molar 86 g / mol consumes 1568 L of air (20% O2) and 180g of water is formed. What is the formula
molecular weight of the substance?
1
Expert's answer
2020-11-23T07:34:28-0500

The combustion reaction is:

2 CxHyOz + (2x+y/2-z) O2 = 2x CO2 + y H2O

Number of moles of organic compound CxHyOz is 172 g / 86 g/mole = 2 mole

Number of moles of O2 consumed is 1568 L * 0.20 / 22.4 L/mole = 14 mole

2x+y/2-z = 14

Number of moles of H2O produced is 180 g / 18 g/mole = 10 mole

y = 10

Therefore:

2x - z = 14 - 5 = 9,

z = 2x - 9.

The molar mass of CxHyOz is 86 g/mole:

12*x + 10 + 16*z = 86.

12*x + 10 + 16*(2x - 9) = 86

x = 5.

Then z = 1

The compound is C5H10O